A transformation T on vector ? of a vector space is defined by T (v ) = A x ? Where the given vector A W and x means the vector product. Find (T2 + A2 T ) ( V ).

Please help me solve this
Answer #1

Trat bad=b-an p = { 142, 1-23 (= {15 n} . T(1+2) = 1–2 T(1-x) =-1-n = 1.1 +E n =-1.1+(1) (Tice & [ - ] Ginen & -P(l) = 4+20 v

answered by: ANURANJAN SARSAM